dddarryl:

how many molecules of carbon monoxide are needed to react with excess iron oxide to produce 11.6 g of iron

justjm:

First you need the equation. \( Fe_2O_3~+~CO→Fe~+~CO_2\) Now, it is not balanced. This is a rather difficult equation to balance because every coefficient on the Carbon influences the oxygen. But you would end up getting the balanced equation as \(Fe_2O_3~+~3CO→2Fe~+~3CO_2\) Now use dimensional analysis. You know that the limiting reagent is the CO, and it's asking you for CO. \(11.6g\ Fe\cdot\frac{1\ mol\ Fe}{55.845g/ mol\ Fe}\cdot\frac{3\ mol\ CO}{2\ mol\ Fe}\cdot\frac{6.023\cdot10^{23}~molecules\ CO}{1\ mol\ CO}~=\) Calculate and you should approach your answer.

dddarryl:

Thank you I was able to complete with this set up

justjm:

NP. Dimensional analysis is your secret trick to any computational work in chemistry or physics. or even economics.